# Iterative Partition Search Variational Quantum Algorithm for Solving Shortest Vector Problem

### v1

@@ III  Iterative Partition Search Algorithm
[…]
5:Pop partition $Y_{i}$ from $S$.
[…]
7:Replace $\bm{b}_{i}$ with $\bm{v}$ and keep $B$ ordered.
8:Push partitions $Y_{i},\cdots,Y_{r}$ onto $S$ in sequence,
[…]
$~Y_{i}$ $=$ $\{(y_{1},\cdots,y_{i-1},1,0,\cdots,0)^{T}\in\mathbb{Z}^{n}\}$ (7)
$~\bm{v}=\sum_{j=1}^{i-1}y_{j}\bm{b}_{j}+\bm{b}_{i}$ (8)
(a)  HEA
(b)  QAOA Ansatz
Figure 1:  Comparison of PQC architectures. (a) HEA: one layer of RY rotations followed by CNOT entanglement and another RY layer. (b) QAOA: alternating objective and mixing layers. For the SVP application, the objective layer is implemented with all-to-all $Z$-interactions, resulting in a deep circuit.
This design, which fixes the coefficient of the $i$-th basis vector to 1, offers several crucial advantages. It reduces qubit requirements by eliminating the encoding for one coefficient. It also simplifies the lattice basis update procedure. Since the $i$-th coefficient is fixed to 1, any solution $\bm{v}$ found in $Y_{i}$ can directly replace $\bm{b}_{i}$ if $\|\bm{v}\|<\|\bm{b}_{i}\|$ while preserving the lattice structure, as formalized in Theorem 1. Most critically, this constrained partitioning simplifies the cost function landscape, mitigating interference from other near-optimal vectors of very similar length. Consequently, the VQA’s search for the true optimum becomes more focused, which significantly increases the probability of convergence and improves the overall solution quality. The efficacy of the 1-tailed search spaces effect is experimentally validated in Section IV.4.
Then, by integrating these focused search steps into an iterative framework, we define the core of our 1-tailed partition strategy. This strategy allows the algorithm to collect multiple distinct candidate solutions across its iterations. The final identification of the shortest vector is then made through a deterministic comparison of these candidates, which enhances the overall probability of success.
Theorem 1
Since the primary focus of our work is not on developing a novel PQC, we adopt a HEA architecture similar to that used in Ref. [25]. This choice, over the more common QAOA ansatz, is motivated by the structure of the SVP Hamiltonian. The SVP Hamiltonian (e.g., Eq. 10) features all-to-all interactions among the qubits. For QAOA ansatz, this inherent connectivity translates into a deep and dense objective layer, as shown in Fig. 1(b), increasing its susceptibility to errors on NISQ devices. In contrast, the HEA architecture, illustrated in Fig. 1(a), offers intense expressivity with a typically shallower circuit depth, making it more suitable for navigating the complex energy landscape of SVP. While HEA can be susceptible to barren plateaus, this risk is substantially mitigated in our framework by the reduced 1-tailed search spaces of each VQA instance and our use of a gradient-free classical optimizer. The efficacy of the PQC effect is experimentally validated in Section IV.2.
For coefficient encoding in partition $Y_{i}$, the $i$-th coefficient is fixed at 1, requiring no qubits. The remaining i1i-1 variable coefficients $y_{r}$ (for $r<i$) are each encoded using a specific number of qubits. For the qubit allocation per variable coefficient, we adopt the approach based on experimental observations and related studies [25, 28, 26]. We find that for an $n$-dimensional lattice, representing each variable coefficient $y_{r}$ with $\lfloor\log n\rfloor$ qubits is typically sufficient for IPSA to identify the SVP solution with high probability. Therefore, an appropriate encoding for these variable coefficients is:
$\hat{y_{r}}=\frac{1}{2}-\sum_{j=0}^{\lfloor\log n\rfloor-1}2^{j-1}\sigma^{z}_{r,j}$ (9)
This encoding represents integers in the range $({-2}^{\lfloor\log n\rfloor-1},2^{\lfloor\log n\rfloor-1}]$. The Hamiltonian for partition $Y_{i}$ is then given by:
$H_{i}^{\prime}=\sum_{k=1}^{n}\left(\sum_{r=1}^{i-1}\hat{y_{r}}b_{r,k}+b_{i,k}\right)^{2}$ (10)
Figure 2:  Qubit requirements of IPSA versus the PSA with different classical preprocessing methods. The plot compares IPSA (blue) with PSA using LLL-reduced bases (red) and HKZ-reduced bases (green). While its qubit scaling is comparable to PSA with HKZ, IPSA avoids an exponential-time HKZ preprocessing routine and maintains a significant advantage over PSA with LLL.
This qubit allocation strategy requires a total of $(n-1)\lfloor\log n\rfloor$ qubits for the largest partition $Y_{n}$ since only n1n-1 coefficients are variable and require qubit representation. When the qubit count of the PSA is set by boundary constraints derived from an LLL-reduced basis, it needs $n(n+1)$ qubits, significantly more than IPSA requires. In the case of using constraints from an HKZ-reduced basis, PSA exhibits a comparable $\mathcal{O}(n\log n)$ scaling to IPSA. However, the HKZ-based boundary analysis relies on a classical preprocessing routine with exponential-time complexity, which is often computationally prohibitive to implement. A visual comparison of qubit requirements is shown in Fig. 2
Although theoretical analyses, such as those boundary constraint methods based on LLL and HKZ basis properties to ensure inclusion of the shortest vector, provide valuable upper bounds and specific allocation schemes [25, 26], empirical validation and heuristic insights remain crucial. These are essential for developing qubit allocation strategies that optimize the trade-off between solution quality and qubit count for practical VQA-based SVP solvers like IPSA, especially when aiming to minimize qubit usage beyond worst-case theoretical guarantees.
IV  Numerical Simulations and Results
In this section, we present a systematic numerical comparison of our proposed IPSA against two representative existing methods: the PSA [31] and the IQOAP [28]. All simulations were performed using the Qiskit framework. In Section IV.1, we describe the two distinct sets of SVP instances generated for this study and detail the complete experimental setup for all algorithms. Subsequent sections provide a quantitative analysis across various metrics, confirming the superiority of IPSA.
IV.1  Instances Generation and Experimental Setup
We generated two distinct sets of SVP instances, hereafter referred to as the Benchmark Set and the LLL-Challenging Set, totaling 800 instances across different dimensions. The Benchmark Set comprises 600 instances (200 for each dimension $n\in\{4,5,6\}$) created using a standard procedure [28, 30]: applying random unimodular transformations to the reduced lattice basis. This approach is known to produce instances with difficulty equivalent to that of random lattices [43, 30]. The LLL-Challenging Set consists of an additional 200 instances engineered for $n=6$. These instances were selected by applying the classical LLL algorithm to a large number of randomly generated lattices and retaining only those for which LLL failed to find the true shortest vector, instead finding a different short vector whose length is, on average, only $\approx 2\%$ longer. The ability to solve such instances critically tests an algorithm’s capacity to identify the true shortest vector in the presence of other very short lattice vectors, a scenario where we hypothesize our proposed algorithm offers a distinct advantage.
For coefficient encoding, both IQOAP and our IPSA employ two qubits per variable coefficient. This choice is consistent with the original IQOAP study [28] and satisfies the $\lfloor\log n\rfloor$ qubit requirement for IPSA in this dimensional range. Since the original PSA proposal [31] does not specify a qubit allocation strategy, we benchmarked several variants, denoted as $k$-PSA, where each coefficient is encoded with $k\in\{3,4,5\}$ qubits. However, the computational cost of classically simulating instances requiring more than 22 qubits was prohibitive. Consequently, we had to exclude the 5-PSA variant for $n=5$ and both the 4-PSA and 5-PSA variants for $n=6$.
To evaluate the choice of PQC within our framework, we also implemented an IPSA variant using the QAOA ansatz (IPSA-QAOA) for comparison against the standard HEA-based IPSA. A direct comparison between different PQCs is complex, involving a trade-off between circuit depth, parameter count, and expressivity. To ensure the QAOA ansatz with sufficient expressivity for solving SVP while maintaining a manageable circuit depth, we configured IPSA-QAOA with $p=4$ layers. In contrast, the standard IPSA-HEA was configured with $p=2$ layers. This configuration establishes a meaningful basis for comparing different PQCs for SVP. The key parameters for all algorithms, including the number of layers $p$ in the PQC, are summarized in Table 1.
Table 1:  Parameter configurations for the simulated algorithms. Here, $n$ is the SVP dimension, and $p$ is the number of layers in the PQC. The settings for IQOAP, including a single-layer QAOA ansatz with constrained parameters, are chosen to be consistent with the original study [28].
Algorithms $n$ Max. Qubits $p$ PQC
IPSA 4 6 2 HEA
5 8
6 10
IPSA-QAOA 4 6 4 QAOA
5 8
6 10
IQOAP 4 8 1 QAOA($\beta=\gamma$)
5 10
6 12
3-PSA 4 11 2 HEA
5 14
6 17
4-PSA 4 15 2 HEA
5 19
5-PSA 4 19 2 HEA
For the classical optimization component of all algorithms, we employed the minimize function from the SciPy.optimize library, selecting Powell’s conjugate direction method as the optimizer. The initial parameters for the rotational gates in the PQC were randomly initialized from a uniform distribution over the interval $0,\pi]$. We adopted the default termination criteria from the SciPy implementation: the optimization process stops when the fractional tolerance in either the parameters (xtol) or the cost function value (ftol) is below $1\times 10^{-4}$. All simulations were performed on the Qiskit framework using the StatevectorEstimator or StatevectorSampler.
@@ IV.1  Instances Generation and Experimental Setup
Figure 3:  Comparative analysis of IPSA-HEA (blue) and IPSA-QAOA (orange) on the Benchmark Set for dimensions $n\in\{4,5,6\}$. The four panels show (a) Success Rate (SR), (b) Approximation Ratio (AR), (c) Total Circuit Depth ($D_{\text{total}}$), and (d) Total CNOT Count ($C_{\text{total}}$). The y-axis values for panels (c) and (d) are presented in units of $10^{6}$. In all plots, the shaded areas represent the interquartile range (25th to 75th percentiles), and the cross markers indicate the median values.
Throughout the following sections, we assess algorithm effectiveness using four key metrics:
@@
    Approximation Ratio (AR): The quality of the found solution for a given instance, defined as $AR=\lambda_{1}(\mathcal{L})/\|v_{\text{alg}}\|$, where $\lambda_{1}(\mathcal{L})$ is the length of the shortest vector and $\|v_{\text{alg}}\|$ is the length of the vector found by the algorithm. An AR value ranges from 0 to 1, with a value closer to 1 indicating a higher-quality solution. The Average Approximation Ratio (AAR) is then calculated for each algorithm and dimension by taking the mean of the AR values over the 200 corresponding instances.
@@
    Total CNOT Count ($C_{\text{total}}$): The cumulative number of two-qubit CNOT gates used across all iterations, $C_{\text{total}}=\sum_{i=1}^{I}c_{i}$. This metric quantifies the primary source of noise and error in many NISQ devices.
IV.2  Comparative Analysis: HEA versus QAOA in IPSA
We begin by empirically validating our choice of PQC architecture. This section involves comparing the relative effectiveness of the IPSA implementation using a 2-layer HEA against the version using a 4-layer QAOA. The comparison was conducted on the 600 instances of the Benchmark Set, and the results are presented in Fig. 3.
The figure shows a clear superiority of the HEA implementation across all dimensions. For instance, at $n=6$, the HEA variant achieved an SR of 0.95, whereas the QAOA variant’s SR was only 0.025. The AAR for IPSA-HEA also remained high and close to 1. Its AAR value for $n=6$ was 0.9835. In contrast, the AAR for IPSA-QAOA degraded markedly as the problem size increased, falling from 0.6254 at $n=4$ to 0.4179 at $n=6$.
@@ IV.2  Comparative Analysis: HEA versus QAOA in IPSA
Figure 4:  Comparative analysis of IPSA (blue) and several existing algorithms, including the $k$-PSA variants (3-PSA in purple, 4-PSA in orange, and 5-PSA in green) and IQOAP (red). The comparison is across dimensions $n\in\{4,5,6\}$ on the Benchmark Set. The panels show (a) Success Rate (SR), (b) Approximation Ratio (AR), (c) Total Circuit Depth ($D_{\text{total}}$), and (d) Total CNOT Count ($C_{\text{total}}$). The y-axis values for (c) and (d) are in units of $10^{5}$. In panels (a) and (b), the brown bars labeled “Initial” represent the metrics of the initial basis. For the data points, the error bars indicate the interquartile range (25th to 75th percentile), and the circular markers denote the median values.
@@
Furthermore, the HEA-based approach was substantially more resource-efficient. At $n=6$, $D_{\text{total}}$ for IPSA-HEA was $0.321\times$106. This result is approximately five times lower than the $1.638\times$106 required by IPSA-QAOA. Its $C_{\text{total}}$ showed an even greater disparity. The value was $0.242\times$106 for IPSA-HEA, nearly ten times lower than the $2.344\times$106 for the QAOA variant. Similar resource advantages for the HEA implementation were observed for $n=4$ and $n=5$.
@@
In summary, these results provide strong empirical evidence that for the SVP instances under consideration within the IPSA framework, the HEA offers a clear and comprehensive advantage over the standard QAOA ansatz, delivering superior solution quality with significantly lower resource consumption. Therefore, for all subsequent experiments presented in this paper, the HEA is adopted as the default PQC for IPSA. While we acknowledge that specific QAOA variants might outperform the standard QAOA ansatz, a detailed investigation of these alternatives is beyond the scope of this work and remains a direction for future research.
IV.3  Comparison with Existing Algorithms

### v2

@@ III.2 Stack-Managed Iterative Process
4: Push partitions $Y_{n},\cdots,Y_{1}$ onto $S$ ($Y_{1}$ at top).
5: while $S$ is not empty do
6:   Pop partition $Y_{i}$ from $S$.
7:   Solve SVP for $Y_{i}$ using HEA, obtaining the result $\bm{v}$.
8:   if $\|\bm{v}\|<\|\bm{b}_{i}\|$ then
9:    Replace $\bm{b}_{i}$ with $\bm{v}$ and keep $B$ ordered.
10:    Push partitions $Y_{i},\cdots,Y_{r}$ onto $S$ in sequence, 11:where $r$ is the position of $\bm{v}$ in $B$.
Since the primary focus of our work is not on deriving new optimal qubit encoding schemes. We follow the strategy demonstrated in the IQOAP framework [zhu2022iterative], which employs a small, fixed bit allocation.
III.3 Parameterized Quantum Circuit Selected
The design of the PQC also plays a role in VQA effectiveness. Standard ansatz like QAOA, while effective for certain optimization problems [ni2024multilevel, wu2025resource, ni2025progressive], typically require deep circuit architectures for SVP due to the all-to-all connectivity inherent in the problem Hamiltonian (e.g., Eq. 6). This limitation motivates the adoption of alternative circuit structures, such as the Hardware-Efficient Ansatz (HEA) [cong2019quantum, henderson2020quanvolutional, liu2021hybrid, hur2022quantum, herrmann2022realizing], which can provide strong expressive power with shallower circuits. Recently, we have also noted other mitigation strategies, such as using Fixed-Angle QAOA variants [boulebnane2024solving, prokop2025heuristic, priestley2025practically] or leveraging classical post-processing [hou2025effcient].
On the superconducting quantum processor, this all-to-all connectivity requirement presents a significant challenge during transpilation. The QAOA ansatz, when applied to the problem Hamiltonian, must implement entangling operations such as $\text{R}_{\text{ZZ}}(\theta)$ gates between all pairs of qubits. These logical entangling gates typically decompose into CZ gates. However, current superconducting quantum processors have limited qubit connectivity, with all-to-all connections not physically available. To execute gates between non-adjacent qubits, the transpiler must insert a substantial number of SWAP gates, which themselves also decompose into multiple CZ gates (Fig. 1). This process results in a significant “transpile tax” overhead. The resulting physical circuit depth increases significantly, making the algorithm highly susceptible to gate errors and decoherence within the device’s finite coherence times.
In contrast, the HEA structure is Hamiltonian-agnostic. Its entangling blocks can be designed to align with the native hardware topology. This alignment drastically reduces the required number of SWAP gates during transpilation. Consequently, the physical HEA circuit remains comparatively shallow and is more robust to hardware noise. This stark contrast in compilation outcomes is visualized in Fig. 2, and the quantitative resource comparison is detailed in Table 1. We provide a detailed performance comparison between QAOA and HEA under ideal, noiseless simulation conditions in Appendix A, which further supports our selection. Therefore, we utilize the HEA as the PQC structure for IPSA.
Figure 1:  Decomposition of logical gates required by the QAOA ansatz. An $\text{R}_{\text{ZZ}}(\theta)$ rotation gate is decomposed into a sequence involving two CZ gates. A SWAP gate is decomposed into three CZ gates.
Figure 2:  Visualization of the compilation overhead for different ansatz structures. The logical circuit diagrams are shown alongside their corresponding physical circuits after transpilation for a linear connectivity topology. The all-to-all connectivity requirement of the QAOA ansatz, which is optimized by Qiskit’s transpiler, still results in a deep physical circuit. The HEA, designed to match the linear topology, undergoes minimal change during transpilation.
Table 1:  A comparison of the logical and transpiled 1-layer QAOA ansatz and the 1-layer HEA for a 6-qubit implementation to be run on real hardware in terms of depth and number of entangling gates. We consider both unoptimized and optimized transpilation using Qiskit’s transpiler (optimization levels 0 and 3) before execution on the Baihua superconducting quantum processor via the Quafu platform.
Logical Circuits
Ansatz Depth Entangling Gates
QAOA 12 15
HEA 7 5
Transpiled Circuits
(Basis Gate: H, Rx, Ry, Rz, CZ)
Ansatz Depth CZ Gates
QAOA (unoptimized) 92 51
QAOA (optimized) 64 43
HEA (unoptimized) 13 5
HEA (optimized) 13 5
IV Real Hardware Experiments
In this section, we present the experimental validation of our proposed IPSA. We benchmark its performance against IQOAP and PSA, through execution on the Baihua superconducting quantum processor, which is accessed via the Quafu platform [Quafu]. Detailed specifications of this hardware and its topology are provided in Appendix B. We will first detail the complete experimental setup and subsequently present the comparative results obtained under realistic noise conditions.
IV.1 Experimental Setup
For the experimental validation, we generated a test set of 50 random SVP instances in dimension $n=4$. This dimension was selected as a representative case for hardware execution, given the significant impact of device noise, limited coherence times, and overall execution time constraints on larger-scale problems. These instances were created using the standard procedure of applying random unimodular transformations to LLL-reduced lattice bases [zhu2022iterative, barbera2024finding].
For the qubit allocation, reflecting the qubit-saving characteristic of iterative frameworks, both IPSA and IQOAP encode each variable coefficient using two qubits, whereas PSA utilizes three qubits per coefficient. For the $n=4$ instances, this results in a circuit with 6 qubits for IPSA, 8 for IQOAP, and 11 for PSA. To ensure a fair comparison, the coefficient range for all test instances was restricted to be representable within the 3-qubit allocation used by PSA.
Regarding the PQC structure, both IPSA and PSA employ a 1-layer HEA. The IQOAP implementation uses a 1-layer QAOA with parameters constrained such that $\beta=\gamma$, consistent with its original proposal [zhu2022iterative].
@@ IV.1 Experimental Setup
Figure 3:  Comparative analysis of IPSA (blue), PSA (purple), and IQOAP (red) on the 50 SVP instances at $n=4$, executed on the Baihua superconducting quantum processor. The panels show (a) Success Rate (SR), (b) Approximation Ratio (AR), (c) Total Circuit Depth ($D_{\text{total}}$), and (d) Total CZ Count ($C_{\text{total}}$). For the data points, the error bars indicate the interquartile range (25th to 75th percentile), and the markers denote the median values.
To quantitatively assess algorithm performance, we utilize the following metrics:
@@
    Approximation Ratio (AR) and Average Approximation Ratio (AAR): The quality of the found solution for a given instance, defined as $AR=\lambda_{1}(\mathcal{L})/\|v_{\text{alg}}\|$, where $\lambda_{1}(\mathcal{L})$ is the length of the shortest vector and $\|v_{\text{alg}}\|$ is the length of the vector found by the algorithm. An AR value ranges from 0 to 1, with a value closer to 1 indicating a higher-quality solution. The AAR is then calculated for each algorithm and dimension by taking the mean of the AR values over all instances.
@@
    Total CZ Count ($C_{\text{total}}$): The cumulative number of two-qubit CZ gates used across all iterations, $C_{\text{total}}=\sum_{i=1}^{I}c_{i}$. This metric quantifies the primary source of noise and error in many NISQ devices.
IV.2 Experimental Results and Analysis
We now present the results from the experimental execution on the Baihua superconducting quantum processor, comparing IPSA against PSA and IQOAP. The performance of the three algorithms on the 50 test instances is summarized in Fig. 3.
As shown in Fig. 3(a), IPSA achieves a Success Rate (SR) of 0.84, which is approximately 2.5 times the SR of 0.34 obtained by IQOAP. In contrast, the SR for PSA was significantly lower at 0.06. Regarding solution quality, IPSA also demonstrates strong performance, reaching an Average Approximation Ratio (AAR) of 0.887. This value is roughly 1.8 times higher than the 0.492 achieved by IQOAP and 5 times higher than the 0.178 of PSA, as depicted in Fig. 3(b).
The hardware experiments also highlight the resource efficiency of IPSA. As detailed in Fig. 3(c), the $D_{\text{total}}$ for IPSA was $1.299\times 10^{4}$, representing an 82.7% reduction compared to the $7.505\times 10^{4}$ required by IQOAP. This significant difference is a direct consequence of the heavy “transpile tax” on real hardware, as discussed in Sec. III.3. A similar trend is observed for $C_{\text{total}}$ in Fig. 3(d), where IPSA uses 91.0% fewer CZ gates compared to the $5.042\times 10^{4}$ count of IQOAP.
In summary, the experimental results on real quantum hardware demonstrate that IPSA achieves a superior balance between solution quality and resource consumption on current superconducting quantum processors. Compared to PSA, IPSA achieves a 14-fold increase in SR with less than double the total circuit depth. Conversely, compared to IQOAP, IPSA reduces $D_{\text{total}}$ by approximately 83% while achieving approximately 2.5 times the SR of IQOAP. Since $D_{\text{total}}$ reflects the cumulative quantum circuit execution time to some extent, this reduction suggests a significant advantage for IPSA in overall runtime efficiency when compared to IQOAP. Simultaneously, the high SR relative to PSA indicates that the dynamic, stack-managed IPSA is more resilient to local minima than static search methods, even under realistic noise conditions. These outcomes suggest that by addressing the structural inefficiencies of its predecessors, IPSA provides a more practical path for applying iterative VQAs to combinatorial optimization problems on near-term quantum devices.
In addition, these real hardware results are consistent with findings from extensive noiseless numerical simulations presented in Appendix C, which demonstrate similar performance advantages for IPSA on larger and more challenging sets of SVP instances.
V  Conclusion
In this work, we proposed IPSA, a “second-generation” iterative VQA-based SVP solver that advances the frameworks of IQOAP and PSA. It is designed to overcome their fundamental drawbacks, including ineffective iterations, oversized search spaces, and low circuit efficiency. Its core features, the 1-tailed search spaces and a dynamic, stack-managed iterative process, are the key components for achieving this goal. Although PSA and IQOAP may not be the best approach for solving SVP with VQA today, by systematically addressing these known limitations, the refined IPSA presented in this work emerges as a competitive approach. Furthermore, IPSA can be employed as a subroutine within algorithms like as block Korkin-Zolotarev or applied to problems such as learning with errors, providing an efficient and reliable core component for these complex tasks.
Acknowledgements.
This work is supported by National Natural Science Foundation of China (Grant Nos. U25B2014, 62372048, 62272056, 62371069), and National Key Laboratory of Security Communication Foundation (2025, 6142103042503). We acknowledge the support of Quafu Quantum Cloud Computing Platform from Beijing Academy of Quantum Information Sciences (https://quafu-sqc.baqis.ac.cn/).
Appendix A Comparative Analysis of HEA and QAOA Ansatz under Ideal Simulation
This appendix provides an empirical validation for the selection of the HEA over the QAOA ansatz. The comparisons presented here are conducted using ideal, noiseless numerical simulations which assume all-to-all qubit connectivity. The definitions for the performance metrics used in this section are provided in Sec. IV.1.
Acknowledging the different expressive power of the two structures, we compare the relative effectiveness of the IPSA implementation using a 1-layer HEA against the version using a 4-layer QAOA. The comparison was conducted on a set of 600 SVP instances (200 for each dimension $n\in\{4,5,6\}$). These instances were created using a standard procedure: applying random unimodular transformations to the reduced lattice basis [zhu2022iterative, barbera2024finding]. The results are presented in Fig. 4
@@ Appendix A Comparative Analysis of HEA and QAOA Ansatz under Ideal Simulation
Figure 4:  Comparative analysis of IPSA-HEA (blue) and IPSA-QAOA (orange) on the set of 600 SVP instances for dimensions $n\in\{4,5,6\}$. The four panels show (a) Success Rate (SR), (b) Approximation Ratio (AR), (c) Total Circuit Depth ($D_{\text{total}}$), and (d) Total CZ Count ($C_{\text{total}}$). The y-axis values for panels (c) and (d) are presented in units of $10^{6}$. In all plots, the shaded areas represent the interquartile range (25th to 75th percentiles), and the cross markers indicate the median values.
The figure shows a clear performance difference when using the HEA implementation across all dimensions. For instance, at $n=6$, the HEA variant achieved an SR of 0.95, whereas the QAOA variant’s SR was only 0.025. The AAR for IPSA-HEA also remained high and close to 1. Its AAR value for $n=6$ was 0.9835. In contrast, the AAR for IPSA-QAOA decreased as the problem size increased, falling from 0.6254 at $n=4$ to 0.4179 at $n=6$.
@@
Furthermore, the HEA-based approach was substantially more resource-efficient in simulation. At $n=6$, $D_{\text{total}}$ for IPSA-HEA was $0.321\times 10^{6}$. This result is approximately five times lower than the $1.638\times 10^{6}$ required by IPSA-QAOA. Its $C_{\text{total}}$ showed an even greater difference. The value was $0.242\times 10^{6}$ for IPSA-HEA, almost ten times lower than the $2.344\times 10^{6}$ for the QAOA variant. Similar resource advantages for the HEA implementation were observed for $n=4$ and $n=5$.
We note that while the HEA variant’s total cost is substantially lower in the tested range, its scaling trend appears steeper than that of the QAOA variant. This can be viewed as the necessary computational trade-off for an ansatz capable of achieving a high success rate, whereas the 4-layer QAOA variant fails to find a solution. We acknowledge that a detailed analysis of the precise asymptotic scaling for both ansatzes is a subject for future investigation.
@@
In summary, these simulation results provide strong empirical evidence that for the SVP instances under consideration within the IPSA, the HEA offers a clear advantage over the standard QAOA ansatz. It delivers superior solution quality with significantly lower resource consumption under ideal conditions. This results complements the hardware-compilation analysis in the main text and supports the adoption of HEA as the PQC for IPSA.
Appendix B Hardware Specifications
In this appendix, we provide additional details about the experimental hardware used in the main text. The experiments were executed on the Baihua superconducting quantum processor, accessible via the Quafu Superconducting Quantum Computing cloud [Quafu]. This processor features 156 superconducting qubits. The qubit connectivity map for two-qubit entangling gates is shown in Fig. 5. The median calibration data for the device, retrieved at the time of the experiments, is summarized in Table 2.
Figure 5:  The qubit connectivity topology of the Baihua superconducting quantum processor. Blue circles indicate available qubits, while light gray circles represent unavailable qubits. Blue arrows denote available connections for two-qubit gates.
Table 2: Median calibration data for the Baihua superconducting quantum processor, representative of the device performance during the experimental period.
Parameter Value
Experimental period October 23 – December 22, 2025
Basis gates H, Rx, Ry, Rz, CZ
[…]
Median 1Q error 7.6e-4
Median 2Q error 1.4e-2
Appendix C Extended Numerical Simulations
In this appendix, we present comprehensive numerical simulations that extend the experimental validation provided in the main text. Specifically, we compare IPSA against IQOAP and multiple variants of PSA (3-, 4-, and 5-qubit encodings) on a larger-scale Benchmark Set. Furthermore, we evaluate algorithm performance on a specialized “LLL-Challenging Set” designed to test robustness against strong suboptimal solutions.
C.1 Instances Generation and Setup
The noiseless numerical simulations utilize two distinct sets of SVP instances, totaling 800 instances.
  • Benchmark Set: Comprises 600 instances (200 each for dimension $n\in\{4,5,6\}$). These were generated by applying random unimodular transformations to reduced lattice bases. Crucially, to ensure a rigorous evaluation of search capabilities, we employed unimodular matrices with larger entries than main-text instances, thereby creating more challenging lattice bases with highly skewed geometries.
  • LLL-Challenging Set: Consists of 200 instances at $n=6$. These instances were specifically engineered by selecting random lattices where the classical LLL algorithm fails to find the true shortest vector, instead converging to a suboptimal vector that is, on average, only $\approx 2\%$ longer. This set critically tests an algorithm’s ability to discriminate between the global optimum and close local minima.
We expanded the comparative analysis to include variants of the PSA algorithm with different qubit allocations, denoted as $k$-PSA, where each coefficient is encoded using $k\in\{3,4,5\}$ qubits. Due to the prohibitive classical simulation costs for circuits exceeding 22 qubits, the 5-PSA variant was excluded for $n=5$, and both 4-PSA and 5-PSA were excluded for $n=6$. The key parameters for all simulated algorithms are summarized in Table 3.
Table 3: Parameter configurations for the algorithms evaluated in the noiseless numerical simulations. The settings for IQOAP are consistent with its original study [zhu2022iterative].
Algorithms SVP dimension Max. Qubits PQC
IPSA 4 6 HEA
5 8
6 10
IQOAP 4 8 QAOA($\beta=\gamma$)
5 10
6 12
3-PSA 4 11 HEA
5 14
6 17
4-PSA 4 15 HEA
5 19
5-PSA 4 19 HEA
For the classical optimization component of all algorithms, we employed the minimize function from the SciPy.optimize library, selecting Powell’s conjugate direction method as the optimizer. The initial parameters for the rotational gates in the PQC were randomly initialized from a uniform distribution over the interval $0,\pi]$. We adopted the default termination criteria from the SciPy implementation: the optimization process stops when the fractional tolerance in either the parameters (xtol) or the cost function value (ftol) is below $1\times 10^{-4}$. All simulations were performed on the Qiskit framework using the StatevectorEstimator or StatevectorSampler. The definitions for the performance metrics used in this section are provided in Sec. IV.1, as same as the main text uses.
C.2 Results on the Benchmark Set